Find the principal needed now to get the given amount, that is, find the present value.
To get $200 after 3 years at 10% compounded monthly
The present value of $200 is $?

Answer by mathmate(429)

Question:
Find the principal needed now to get the given amount, that is, find the present value.
To get $200 after 3 years at 10% compounded monthly

Solution:
F=200
APR=10%
i=0.1/12 [monthly interest]
n=3*12=36 [months]

Use present value formula:
P=F/(1+i)^n
=200/(1+0.1/12)^36
=200/1.34818
=$148.35 [present value required to accumulate $200 after 3 years at APR 10% compounded monthly]